y = x - 2
(0, ), ( ,0), (5, )
Please help me to solve! thanks

Answer by funmath(2933)

Points come in the order (x,y).
y = x - 2
(0, ), ( ,0), (5, )
For (0, ), let x=0 and solve for y:
y=0-2
y=-2
The ordered pair is (0,-2)
:
For ( ,0), let y=0 and solve for x:
0=x-2
2=x-2+2
2=x
The ordered pair is (2,0)
:
For (5, ), let x=5 and solve for y
y=5-2
y=3
The ordered pair is (5,3)
Happy Calculating!!!